OpenStudy (anonymous):

Find three consecutive odd integers such that the sum of the largest and twice the smallest is 25. If x represents the smallest integer, then which equation could be used to solve the problem?

OpenStudy (anonymous):

selecting three consecutive odd integers would make sense if u select like this 2x+1,2x+3 and 2x+5 bcoz if u select x then how can u justify that it is odd.

OpenStudy (mertsj):

Let x = first odd integer x+2 = 2nd odd integer x+4 = third odd integer x+4+2x=25

OpenStudy (anonymous):

And then combine terms... :-) 3x + 4 = 25